Example 1. Mr. Orsini Makes the Call

Mr. Orsini thinks the stock of Apple (symbol AAPL) will go up in value. On March 19, 2021, he buys a call option for 100 shares of AAPL at a strike price of $410. The current share price is $397.77. The cost of the call option, or premium, is $5 per share, for a total of $500. The call option will expire on June 30, 2021.

If the AAPL stock goes up to $420 and the premium goes to $9 before the expiration date, Mr. Orsini has an “in the money” option, which means that the stock has increased in value to more than the strike price. It’s time for Mr. Orsini, who is now very proud of his shrewd investing tactics, to decide how to cash in. He has a few choices: he can exercise the option and purchase the stock, allowing him to buy stock worth $42,000 (100 x $420) for $4,100 (100 x $410). How much would he make in profits? SHOW YOUR WORK!!

If he sold the option, how much would he make? SHOW YOUR WORK!!



1
Expert's answer
2021-03-24T20:11:51-0400
"solution"

100 shares at "\$410"

current share price "\$397.77"

call option, 100 shares at each"\$5 ..i.e\ 100\times 5=\$500"

current cost of shares will be :

"397.77\times 100=39,777"


profit: if purchased shares

"41,000-39,777=\$1,223 \\
total\ profit \ \$1223"


selling option :

"\$420\times 100=\$42,000\\
\$42000-\$39777=\$2223"